# The Atacama Cosmology Telescope: The Development of Machine Learning Tools for Detecting Millimeter Sources in Timestream Pre-processing

### v1

@@ 6 Stellar Flare Simulations
To test the classification algorithm on a point source with known position and amplitude, we inject simulated signals of stellar flares into real TODs. This allows us to test our classification pipeline on timestreams that contain real noise plus the addition of signals from simulated sources. We then use the known pointing matrix, which describes the sky coordinates of the telescope pointing as a function of time, to project an instrumental beam profile, as is appropriate for a point source, into the TOD timestreams. This process projects the 2D spatial beam into the timestreams for all detectors that scan over the simulated source.
@@
where is the initial amplitude of the flare, is the half-life of the decaying flare, and is the start time of the flare, which we select to be when the source enters the focal plane for the first time in the TOD. Stellar flares can display a variety of behaviors across the electromagnetic spectrum and can be fit with different profiles. For instance, MacGregor et al. 32 fit two millimeter flares detected by the Atacama Large Millimeter/submillimeter Array 50 ALMA with Gaussian profiles. Salter et al. 42, on the other hand, used an exponential curve to fit the decaying millimeter flux of the equal-mass, highly eccentric, close binary system DQ Tau observed with the Institute for Radio Astronomy in the Millimetre Range Plateau de Bure Interferometer 21 IRAM PdBI, the Combined Array for Research in Millimeter-wave Astronomy 11 CARMA, and the Submillimeter Array 27 SMA. In this paper, we opt for a simple model in which an instantaneous jump in flux is followed by an exponential decrease.
Figure 11 shows an example of a stellar flare injected into a TOD. The simulations are run for sources with an initial amplitude of 12,000 and mJy and flare profiles with half-lives of seconds. These amplitudes were chosen as this specific TOD has an average noise level of mJy, and we wanted to test a case where most detectors would be well above the cut-off and a case closer to the limit. These half-lives were chosen to cover a wide range of flare durations. The shortest ( s) half-life reproduces a situation where the flare decays significantly within a single scan, whereas the longest ( s) half-life was chosen to maximize the number of detections within the 11-minute TOD. The TODs are loaded in raw data acquisition units (DAU) then converted into mJy, at which point the source signal is injected. The TOD is then converted to K to be compatible with TOD loading and cuts packages, at which point it can be run through our pipeline.
@@ 7.1 Classification Results
We tested our trained forest on a set of real data to determine how effective it is at classifying true snippets, all labeled as outlined in Section 5.1. Our training set includes 2580 glitches which consist of 272 PS, 167 PS+, 1119 CR and 1,022 EG objects respectively. Of these glitches, 2280 are from 2019, 90 GHz from PA5, and the remaining 300 are from 2021, 90 GHz and 150 GHz from PA4 and PA5. For testing, we had 3865 glitches: 282 PS, 207 PS+, 1672 CR, and 1704 EG. This consisted of 760 glitches from 2019, 90 GHz, and PA5 as well as 3,105 from 2021, 90 GHz and 150 GHz from PA4 and PA5. Figure 12 shows the resulting confusion matrix, where each value indicates the decimal percentage of each glitch category on the -axis being predicted as the category shown on the -axis. For perfect performance, the resulting confusion matrix would be an identity matrix. Our overall accuracy is 90%, and notably, we achieved a 94% accuracy for PS objects. For CR objects, are labeled EG and of the EG objects are being labeled as CR. This is expected as some electronic glitches, such as large-area heat-ups of the focal plane, look similar to cosmic rays and vice versa.
To test the performance of the classifier, we also examined the precision, recall, and F-1 score (a combination of precision and recall) for each of the classification categories. As illustrated in Equation 5, precision is the proportion of glitches predicted to be within a category that were predicted correctly, while recall measures how many glitches of a certain category were correctly predicted. A high precision score therefore indicates a low false positive rate (few objects incorrectly identified), whereas a high recall indicates a low false negative rate (few objects missed). The F-1 score is a harmonic mean of the precision and recall of the forest. They are given by:
where tp is the number of true positives, fp is the number of false positives, fn is the number of false negatives, is the precision, is the recall, and F-1 is the F-1 score. The results of our test set can be seen in Table 1. All the F-1 scores across the four categories are greater than 89%, which indicates a well-performing model. We consider a well-performing model to be one where the metrics discussed above are ; however, we also favor as high a value of precision as possible, as false detections of point sources that are left in the TODs will have a negative impact on the maps.
@@
We consider a classification of an individual object with a true underlying type to be “high-probability” if it is predicted to be a particular type with a probability, defined as the percentage of decision trees in the forest that predict the glitch to be a given type, to be greater than 70%. After filtering for high-probability classifications in our test set, we are left with six PS+ objects that were mistakenly classified as PS objects, and examine these individually to understand the misclassifications. The TODs and focal plane positions for these snippets are shown in Figure 13. In these TODs, another glitch appears alongside a point source, but the detectors affected by the other glitch are located near to the PS in the focal plane. This resulted in the “Normalized Maximum within 0.1” statistic for all cases to be equal to one. Similarly, the “Normalized Maximum” values for these objects range from 0.27 to 0.51 and the “ and Extent Ratios” range from 0.06 to 0.24, which is consistent with values observed from objects with the PS designation rather than PS+. As described in Section 4, the three statistics mentioned above are usually the only ones that separate cleanly between the PS and the PS+ categories, while the other statistics used for classification are consistent with expected values for both types. In this case, the fact that the statistics we input as the features of our algorithm suggest a PS categorization (as is shown in Figure 2) is the dominant reason for the misclassification. We repeat that these three statistics are important for the classification results, as can be seen in Figure 7. Future work should be done to determine another statistic to aid classification in these instances. One possibility would be a statistic that identifies spikes in the signal that are beam-shaped as we would expect from a source but not from a cosmic ray or electronic glitch. Note that human bias also affects the results, as our training and testing labels are from human labeling.
@@ 7.2 Improved Transient Search Maps
In addition to removing glitches from the TODs prior to map-making, our algorithm is also able to classify astrophysically interesting bright sources from the TODs and generate a list of detection times and locations of these sources. This represents a complementary method to map-based transient search algorithms e.g., those presented in 12. For map-based transient searches, “depth-1” maps, or maps from a single array and frequency in which each point in the map drifts only once through the focal plane—roughly speaking, a single observation session—are used. Biermann et al. 12 provides details on the ACT depth-1 maps and the results of blind transient searches on these maps. Current map-based searches have focused on finding faint sources outside the Galactic plane. This is partially due to the limitations of identifying many sources that are spatially close together in the maps, as occurs within the Galactic plane. Because our algorithm processes each TOD snippet individually, we do not, however, encounter this limitation to the same extent as map-based searches.
@@
To demonstrate that our algorithm can prevent genuine sources from being removed from depth-1 maps, we modify the ACT cuts software to not remove any source which has a probability of being a PS greater than 70%, as determined by the random forest, and compare the results to what is obtained with the original cuts software, doing so on a series of stellar flare simulations described in Section 6. The steps of this modified algorithm can be seen in Figure 14, where the existing portions of the pipeline that we modified are shown in purple and our new glitch classification algorithm is shown in teal. In order to demonstrate the efficacy of our method, we generate a bright, relatively stable source that stays at by injecting a stellar flare with a peak amplitude of 12,000 mJy and a half-life of 5000 s. The results can be seen in Figure 15. We note that the original algorithm removes the source from the map, whereas our modified algorithm restores the source. A ringing pattern along the scan direction through the source is visible in this map cut-out. This is a form of subpixel model error bias (related to regression dilution) caused by treating the sky as being stepwise constant in each pixel despite the true signal changing smoothly, and is unrelated to our cuts implementation 38. Additionally, as we are only producing maps for the purposes of source detection, there could be a lack of convergence in the maps which could also produce some of this affect.
@@
To test the ability of our algorithm to identify sources with different properties, we vary the initial peak amplitude and half-life of the simulation. Recall that the cuts algorithm identifies glitches by locating spikes in the TOD with an ; the TOD used in this simulation has an average noise of mJy, though individual detectors can have more or less noise than this value. We thus start our exploration with an amplitude of 12,000 mJy to ensure that we are comfortably above the threshold. The top row and bottom row per amplitude of Figure 16 shows the maps with each of the original and modified cuts pipelines, respectively. For this large amplitude, the source was at least partially cut out of the maps using the original pipeline, but completely restored with the new pipeline for all simulated source half-lives except 0.5 s, which was classified as a PS, but only with a probability of 70% (which is at the threshold of what we consider a high-probability detection). In the case of the 6000 mJy amplitude source, it is above the noise value for most, but not all, of the detectors. In this case, the source is classified as a PS for and 5 s but only with a probability of 68%, which is below our detection threshold. The source appears in the map made with the original algorithm and also in our modified algorithm for and 250 s and the center of the source is fully restored for and 750 s which is necessary for map-based analysis of the source properties. The reason the source can appear in the original maps even though some detectors are cut is because a portion of the detectors see the source below the threshold, while the ACT cuts algorithm requires that the source has an for a detector to be cut. There is some utility, however, to being able to identify a source that is only partially detected/cut, as one can measure the amplitude variations between snippets to aid classification of the source when it is later detected in a map-based search. Partial identification of sources also generates a list of times and locations of observed sources that are useful for the map-based search methods described above. Note that we chose the probability threshold in order to reduce the number of false positives for map-making, but for transients searches one could lower this threshold and the source would have been recovered for both amplitudes with a half-life of 0.5 s.
@@
Finally, we tested our algorithm for an amplitude of 3000 mJy, which is similar to the amplitude of the brightest transient found in ACT ( mJy; Biermann et al. 12, Li et al. 31). We note, however, that previous transient searches have excluded the galactic plane, where we have found many sources with higher amplitudes that we have successfully identified and restored with the new algorithm. The injected source was only identified for and 750 s with all other values of half-life resulting in the source remaining below the detection threshold. We would like to understand if this failure to correctly classify the injected source is due to a limitation of our classification performance or a result of the earlier step in the process where the snippet is flagged by the cuts algorithm.
@@
To do so we ran the simulated sources through a modified cuts algorithm that we call the “ideal cuts” algorithm, that flags every detector that observes the source with an amplitude greater than 500 mJy, rather than the usual approach used for ACT analysis which requires the signal to be above the threshold. Given that this is a simulation and hence we know a source is present, we can isolate the “detection” versus the “classification” parts of our analysis, removing any uncertainty in the detector cuts pipeline and focusing only on the performance of the classification algorithm presented in this paper. Employing this procedure (where a source with an amplitude greater than 500 mJy detected by any detector is flagged), we were able for each simulated amplitude to detect the injected sources for all values of the half-lives and initial amplitudes we simulated. We provide more details on this test in Appendix Section A. The failure to detect the source by the detector cuts algorithm is due to the source having an below the threshold. Due to the importance of summary statistics related to the focal plane statistics in our classification algorithm (as is shown in Figure 7), we are able to classify the sources, provided they are detected by our “ideal cuts” algorithm even though the source signal is very faint in the TODs. The cutoff is a known limitation of the cuts algorithm, highlighting the need for future work to create an improved cuts algorithm that would enable our classification algorithm to find lower-amplitude sources. An improved cuts algorithm would enable TOD based classification algorithms, such as the one outlined in this paper, to identify dim sources prior to map-making. This would not only aid in quick identification of possibly interesting sources, but also provide a list of candidate positions and times to search for sources in the depth-1 maps.
@@ 7.3 Point Source Variability
If a source is observed multiple times, as the focal plane scans back and forth in azimuth at constant elevation, we can measure changes in the amplitude of the source on scales of less than a minute for an ACT-like experiment. This has important implications for transient science, where some sources, like flaring stars, can vary on time scales of minutes or seconds, whereas other sources, such as extragalactic transients like GRBs, TDEs and SNs, vary on much longer time scales of days or more. To issue timely and accurate alerts of the detection of such events to the scientific community, we need our code to be able to measure changes in amplitude on similar timescales. Here, we demonstrate the capabilities of our algorithm to retrieve variability information on classified point sources using the stellar flare simulations described in Section 6.
@@
where is the amplitude measured by a detector observing the sky at right ascension/declination position , is the intrinsic amplitude of the source located at position , is the azimuthally-averaged, normalized beam profile of the telescope, and is an offset to account for the baseline of the signal being nonzero. The independent variable in the fit is and the fit parameters are , and . Before performing the fit, we create a copy of the snippet with the peaks excised, and we subtract this copy from the original snippet, leaving only the peaks generated by the point source in the data. This removes any remaining large-scale drift across the length of the snippet, which we find significantly improves the fit. We then discard half of the 0.5 s buffer from each of the beginning and end of the snippet, i.e., the first and last 100 samples to reduce the amount of noise that goes into the fit, as we know that the snippets have a buffer of 200 samples on either side of the source signal. Finally, we obtain initial guesses for the non-linear fit by averaging the peak amplitude values of all detectors, for the initial value of , and the detector positions at their peak values, for ; is initially set to zero. We then run the curve_fit non-linear fitter as implemented by SciPy.
@@
We group snippets that are classified as high-probability point sources within 0.01 of each other (corresponding to the standard error on our point source localization in the timestreams), treating them as the same astronomical object, and then we do the fit described above for each scan of the array across the source. This produces a light curve with one point per scan. Our best-fit light curves to simulated stellar flares injected into real data are shown in Figure 17. The position of the source relative to the scan pattern differs for each TOD, causing the time intervals between the source showing up in successive scans to vary from TOD to TOD. This can be seen in the different segment lengths of the individual best-fit light curves (dashed lines) in Figure 17. For this reason, we linearly interpolated the best-fit light curves into 10 s time bins. The amplitudes were then averaged in each time bin to obtain a mean and standard deviation, shown with the shaded region in Figure 17. We only simulate stellar flares with half-lives above 25 seconds because shorter flares decay too quickly to be detected for more than one scan.
@@
To evaluate the level of confidence with which we can detect short-term variations with this method, we calculate the likelihood ratio (or , assuming a Gaussian likelihood) of our best-fit light curve from a single TOD to a flat line at the average flux value; the latter is the null hypothesis of no variation in the source. The likelihood ratio can be converted to a p-value for the null hypothesis, or equivalently a significance of the measurement of variability.333The conversion of the likelihood ratio to a -value assumes that the former has a distribution, which is true for numerous samples; in our case the number of samples is low, so the values we report should be taken as approximate. For uncertainties in our measurement, we use the standard deviation from fits to 11 different TODs into which the same simulated flare is injected. The results of this analysis are presented in Table 2.
@@
750 | 2.8 | <1 |
500 | 4.2 | 1.7 |
250 | 5.8 | 2.0 |
100 | 10.9 | 1.6 |
25 | 5.3 | <1 |
Although the algorithm’s ability to detect a varying signal is currently limited to very bright objects, due to the threshold of the cuts algorithm (as discussed in the previous section), these results highlight the potential of the algorithm to detect quickly decaying flares, sub-minute flares with at least two observations, in the timestreams. This will be an essential tool in detecting fast varying (sub-minute to minute) transients.

### v2

@@ 6 Stellar Flare Simulations
To test the classification pipeline on a point source with known position and amplitude, we inject simulated signals of stellar flares into real TODs. This allows us to test our classification pipeline on timestreams that contain real noise plus the addition of signals from simulated sources. We then use the known pointing matrix, which describes the sky coordinates of the telescope pointing as a function of time, to project an instrumental beam profile, as is appropriate for a point source, into the TOD timestreams. This process projects the 2D spatial beam into the timestreams for all detectors that scan over the simulated source.
@@
where is the initial amplitude of the flare, is the half-life of the decaying flare, and is the start time of the flare, which we select to be when the source enters the focal plane for the first time in the TOD. Stellar flares can display a variety of behaviors across the electromagnetic spectrum and can be fit with different profiles. For instance, MacGregor et al. 32 fit two millimeter flares detected by the Atacama Large Millimeter/submillimeter Array 49 with Gaussian profiles. Salter et al. 41, on the other hand, used an exponential curve to fit the decaying millimeter flux of the equal-mass, highly eccentric, close binary system DQ Tau observed with the Institute for Radio Astronomy in the Millimetre Range Plateau de Bure Interferometer 21, the Combined Array for Research in Millimeter-wave Astronomy 11, and the Submillimeter Array 27. In this paper, we opt for a simple model in which an instantaneous jump in flux is followed by an exponential decrease.
Figure 11 shows an example of a stellar flare injected into a TOD. The simulations are run for sources with an initial amplitude of 12,000 and mJy and flare profiles with half-lives of seconds. These amplitudes were chosen as this specific TOD has an average noise level of mJy, and we wanted to test a case where most detectors would be well above the cut-off and a case closer to the limit. These half-lives were chosen to cover a wide range of flare durations. The shortest ( s) half-life reproduces a situation where the flare decays significantly within a single scan, whereas the longest ( s) half-life was chosen to maximize the number of detections within the 11-minute TOD. The TODs are loaded in raw data acquisition units then converted into mJy, at which point the source signal is injected. The TOD is then converted to K to be compatible with TOD loading and cuts packages, at which point it can be run through our pipeline.
@@ 7.1 Classification Results
We tested our trained forest on a set of real data to determine how effective it is at classifying true snippets, all labeled as outlined in Section 5.1. Our training set includes 2580 glitches which consist of 272 PS, 167 PS+, 1119 CR and 1022 EG objects respectively. Of these glitches, 2280 are from 2019, 90 GHz from PA5, and the remaining 300 are from 2021, 90 GHz and 150 GHz from PA4 and PA5. For testing, we had 3865 glitches: 282 PS, 207 PS+, 1672 CR, and 1704 EG. This consisted of 760 glitches from 2019, 90 GHz, and PA5 as well as 3105 from 2021, 90 GHz and 150 GHz from PA4 and PA5. Figure 12 shows the resulting confusion matrix. Our overall accuracy is 90%, and notably, we achieved a 94% accuracy for PS objects. For CR objects, were classified as EG and of the EG objects were classified as CR. This is expected as some electronic glitches, such as large-area heat-ups of the focal plane, look similar to cosmic rays and vice versa.
The performance metrics for our test set can be seen in Table 2. All the F-1 scores across the four categories are greater than 89%, which indicates a well-performing model. We consider a well-performing model to be one where the metrics discussed above are ; however, we also favor as high a value of precision as possible, as false detections of point sources that are left in the TODs will have a negative impact on the maps.
We consider a classification of an individual object with a true underlying type to be “high-probability” if it is predicted to be a particular type with a probability, defined as the percentage of decision trees in the forest that predict the glitch to be a given type, to be greater than 70%. After filtering for high-probability classifications in our test set, we are left with six PS+ objects that were mistakenly classified as PS objects, and examine these individually to understand the misclassifications. The TODs and focal plane positions for these snippets are shown in Figure 13. In these TODs, another glitch appears alongside a point source, but the detectors affected by the other glitch are located near to the PS in the focal plane. This resulted in the normalized maximum within 0.1 statistic for all cases to be equal to one. Similarly, the normalized maximum values for these objects range from 0.27 to 0.51 and the and extent ratios range from 0.06 to 0.24, which is consistent with values observed from objects with the PS designation rather than PS+. As described in Section 4, the three statistics mentioned above are usually the only ones that separate cleanly between the PS and the PS+ categories, while the other statistics used for classification are consistent with expected values for both types. In this case, the fact that the statistics we input as the features of the random forest suggest a PS categorization (as is shown in Figure 2) is the dominant reason for the misclassification. We repeat that these three statistics are important for the classification results, as can be seen in Figure 10. Future work should be done to determine another statistic to aid classification in these instances. One possibility would be a statistic that identifies spikes in the signal that are beam-shaped as we would expect from a source but not from a cosmic ray or electronic glitch. Note that human bias also affects the results, as our training and testing labels are from human labeling.
@@ 7.2 Improved Transient Search Maps
In addition to removing glitches from the TODs prior to map-making, our classification pipeline is also able to classify astrophysically interesting bright sources from the TODs and generate a list of detection times and locations of these sources. This represents a complementary method to map-based transient search algorithms e.g., those presented in 12. For map-based transient searches, “depth-1” maps, or maps from a single array and frequency in which each point in the map drifts only once through the focal plane—roughly speaking, a single observation session—are used. Biermann et al. 12 provides details on the ACT depth-1 maps and the results of blind transient searches on these maps. Current map-based searches have focused on finding faint sources outside the Galactic plane. This is partially due to the limitations of identifying many sources that are spatially close together in the maps, as occurs within the Galactic plane. Because our classification pipeline processes each TOD snippet individually, we do not, however, encounter this limitation to the same extent as map-based searches.
@@
To demonstrate that our classification pipeline can prevent genuine sources from being removed from depth-1 maps, we modify the ACT cuts software to not remove any source which has a probability of being a PS greater than 70%, as determined by the random forest, and compare the results to what is obtained with the original cuts software, doing so on a series of stellar flare simulations described in Section 6. The steps of this modified pipeline can be seen in Figure 14, where the existing portions of the pipeline that we modified are shown in purple and our new glitch classification pipeline is shown in teal. In order to demonstrate the efficacy of our method, we generate a bright, relatively stable source that stays at by injecting a stellar flare with a peak amplitude of 12,000 mJy and a half-life of 5000 s. The results can be seen in Figure 15. We note that the original pipeline removes the source from the map, whereas our modified pipeline restores the source. A ringing pattern along the scan direction through the source is visible in this map cut-out. This is a form of subpixel model error bias (related to regression dilution) caused by treating the sky as being stepwise constant in each pixel despite the true signal changing smoothly, and is unrelated to our cuts implementation 38. Additionally, as we are only producing maps for the purposes of source detection, there could be a lack of convergence in the maps which could also produce some of this affect.
@@
To test the ability of our pipeline to identify sources with different properties, we vary the initial peak amplitude and half-life of the simulation. Recall that the cuts algorithm identifies glitches by locating spikes in the TOD with an ; the TOD used in this simulation has an average noise of mJy, though individual detectors can have more or less noise than this value. We thus start our exploration with an amplitude of 12,000 mJy to ensure that we are comfortably above the threshold and 6,000 mJy to have a fair number of detectors around the SNR threshold. The top and bottom rows per amplitude of Figure 16 show maps with each of the original and modified cuts pipelines, respectively. For the 12,000 mJy amplitude source, the source was at least partially cut out of the maps using the original pipeline, but completely restored with the new pipeline for all simulated source half-lives except 0.5 s, which was classified as a PS, but only with a probability of 70% (which is at the threshold of what we consider a high-probability detection). In the case of the 6000 mJy amplitude source, it is above the noise value for most, but not all, of the detectors. In this case, the source is classified as a PS for and 5 s but only with a probability of 68%, which is below our detection threshold. The source appears in the map made with the original pipeline and also in our modified pipeline for and 250 s and the center of the source is fully restored for and 750 s which is necessary for map-based analysis of the source properties. The reason the source can appear in the original maps even though some detectors are cut is because a portion of the detectors see the source below the threshold, while the ACT cuts algorithm requires that the source has an for a detector to be cut. There is some utility, however, to being able to identify a source that is only partially detected/cut, as one can measure the amplitude variations between snippets to aid classification of the source when it is later detected in a map-based search. Partial identification of sources also generates a list of times and locations of observed sources that are useful for the map-based search methods described above. Note that we chose the probability threshold in order to reduce the number of false positives for map-making, but for transients searches one could lower this threshold and the source would have been recovered for both amplitudes with a half-life of 0.5 s.
@@
Finally, we tested our pipeline for an amplitude of 3000 mJy, which is similar to the amplitude of the brightest transient found in ACT ( mJy; Biermann et al. 12, Li et al. 31). We note, however, that previous transient searches have excluded the galactic plane, where we have found many sources with higher amplitudes that we have successfully identified and restored with the new classification pipeline. The injected source was only identified for and 750 s with all other values of half-life resulting in the source remaining below the detection threshold. We would like to understand if this failure to correctly classify the injected source is due to a limitation of our classification performance or a result of the earlier step in the process where the snippet is flagged by the cuts algorithm.
@@
To do so we ran the simulated sources through a modified cuts algorithm that we call the “ideal cuts” algorithm, that flags every detector that observes the source with an amplitude greater than 500 mJy, rather than the usual approach used for ACT analysis which requires the signal to be above the threshold. Given that this is a simulation and hence we know a source is present, we can isolate the “detection” versus the “classification” parts of our analysis, removing any uncertainty in the detector cuts pipeline and focusing only on the performance of the classification pipeline presented in this paper. Employing this procedure, we were able for each simulated amplitude to detect the injected sources for all values of the half-lives and initial amplitudes we simulated. We provide more details on this test in Appendix Section A. The failure to detect the source by the detector cuts algorithm is due to the source having an below the threshold. Due to the importance of summary statistics related to the focal plane statistics in our classification pipeline (as is shown in Figure 10), we are able to classify the sources, provided they are detected by our “ideal cuts” algorithm even though the source signal is very faint in the TODs. The cutoff is a known limitation of the cuts algorithm, highlighting the need for future work to create an improved cuts algorithm that would enable our classification pipeline to find lower-amplitude sources. An improved cuts algorithm would enable TOD based classification pipelines, such as the one outlined in this paper, to identify dim sources prior to map-making. This would not only aid in quick identification of possibly interesting sources, but also provide a list of candidate positions and times to search for sources in the depth-1 maps.
@@ 7.3 Point Source Variability
If a source is observed multiple times, as the focal plane scans back and forth in azimuth at constant elevation, we can measure changes in the amplitude of the source on scales of less than a minute for an ACT-like experiment. This has important implications for transient science, where some sources, like flaring stars, can vary on time scales of minutes or seconds, whereas other sources, such as extragalactic transients like GRBs, TDEs and SNs, vary on much longer time scales of days or more. To issue timely and accurate alerts of the detection of such events to the scientific community, we need our code to be able to measure changes in amplitude on similar timescales. Here, we demonstrate the capabilities of our pipeline to retrieve variability information on classified point sources using the stellar flare simulations described in Section 6.
@@
where is the amplitude measured by a detector observing the sky at R.A./decl position , is the intrinsic amplitude of the source located at position , is the azimuthally-averaged, normalized beam profile of the telescope, and is an offset to account for the baseline of the signal being nonzero. The independent variable in the fit is and the fit parameters are , and . Before performing the fit, we create a copy of the snippet with the peaks excised, and we subtract this copy from the original snippet, leaving only the peaks generated by the point source in the data. This removes any remaining large-scale drift across the length of the snippet, which we find significantly improves the fit. We then discard half of the 0.5 s buffer from each of the beginning and end of the snippet, i.e., the first and last 100 samples to reduce the amount of noise that goes into the fit, as we know that the snippets have a buffer of 200 samples on either side of the source signal. Finally, we obtain initial guesses for the non-linear fit by averaging the peak amplitude values of all detectors, for the initial value of , and the detector positions at their peak values, for ; is initially set to zero. We then run the curve_fit non-linear fitter as implemented by SciPy.
@@
We group snippets that are classified as high-probability point sources within 0.01 of each other (corresponding to the standard error on our point source localization in the timestreams), treating them as the same astronomical object, and then we do the fit described above for each scan of the array across the source. This produces a light curve with one point per scan. Our best-fit light curves to simulated stellar flares injected into real data are shown in Figure 17. The position of the source relative to the scan pattern differs for each TOD, causing the time intervals between the source showing up in successive scans to vary from TOD to TOD. This can be seen in the different segment lengths of the individual best-fit light curves (dashed lines) in Figure 17. For this reason, we linearly interpolated the best-fit light curves into 10 s time bins. The amplitudes were then averaged in each time bin to obtain a mean and standard deviation, shown with the shaded region in Figure 17. We only simulate stellar flares with half-lives above 25 s because shorter flares decay too quickly to be detected for more than one scan.
@@
To evaluate the level of confidence with which we can detect short-term variations with this method, we calculate the likelihood ratio (or , assuming a Gaussian likelihood) of our best-fit light curve from a single TOD to a flat line at the average flux value; the latter is the null hypothesis of no variation in the source. The likelihood ratio can be converted to a p-value for the null hypothesis, or equivalently a significance of the measurement of variability.555The conversion of the likelihood ratio to a -value assumes that the former has a distribution, which is true for numerous samples; in our case the number of samples is low, so the values we report should be taken as approximate. For uncertainties in our measurement, we use the standard deviation from fits to 11 different TODs into which the same simulated flare is injected. The results of this analysis are presented in Table 3.
@@
| 750 | 2.8 | <1 |
| 500 | 4.2 | 1.7 |
| 250 | 5.8 | 2.0 |
| 100 | 10.9 | 1.6 |
| 25 | 5.3 | <1 |
Although the pipeline’s ability to detect a varying signal is currently limited to very bright objects, due to the threshold of the cuts algorithm (as discussed in the previous section), these results highlight the potential of the classification pipeline to detect quickly decaying flares, sub-minute flares with at least two observations, in the timestreams. This will be an essential tool in detecting fast varying (sub-minute to minute) transients.